OpenStudy (anonymous):

Solve the system by the method of substitution. y = 3x + 8 y = 2x + 7 A. (1, 9) B. (–1, 5) C. (1, 11) D. (11, 1)

jimthompson5910 (jim_thompson5910):

y = 2x + 7 3x+8 = 2x + 7 ... replace y with 3x+8 (since y = 3x+8) 3x-2x = 7-8 x = ???

OpenStudy (anonymous):

-1

jimthompson5910 (jim_thompson5910):

if x = -1, then y = 3x + 8 y = 3*(-1) + 8 y = ??

OpenStudy (anonymous):

5

jimthompson5910 (jim_thompson5910):

correct, so (x,y) = (-1, 5)